# UI-AGILE: Advancing GUI Agents with Effective Reinforcement Learning and Precise Inference-Time Grounding

### v3

@@ Abstract
The emergence of Multimodal Large Language Models (MLLMs) has driven significant advances in Graphical User Interface (GUI) agent capabilities. Nevertheless, existing GUI agent training and inference techniques still suffer from a dilemma for reasoning designs, ineffective reward, and visual noise. To address these issues, we introduce UI-AGILE for enhancing GUI agents at both training and inference. For training, we propose a suite of improvements to the Supervised Fine-Tuning (SFT) process: 1) a continuous reward function to incentivize high-precision grounding; 2) a “Simple Thinking” reward to balance planning with speed and grounding accuracy; and 3) a cropping-based resampling strategy to mitigate the sparse reward problem and improve learning on complex tasks. For inference, we present decomposed grounding with selection to dramatically improve grounding accuracy on high-resolution displays by breaking the image into smaller, manageable parts. Experiments show that UI-AGILE achieves the state-of-the-art grounding performance on two benchmarks ScreenSpot-Pro and ScreenSpot-v2 while it also exhibits strong general agent capabilities. For instance, using both our training and inference enhancement methods brings 23% grounding accuracy improvement over the best baseline on ScreenSpot-Pro. We provide the code in https://github.com/KDEGroup/UI-AGILE.
@@ 1 Introduction
Driven by the growing capabilities of Multimodal Large Language Models (MLLMs) in image understanding (Bai et al. 2025), Graphical User Interface (GUI) agents, which execute tasks by understanding screenshots and user instructions, are advancing rapidly (Zhang et al. 2025).
Prior methods for GUI agents mostly rely on Supervised Fine-Tuning (SFT), requiring a large amount of human-annotated or synthesized data for teaching agents how to plan its actions and grounding (Gou et al. 2025; Wu et al. 2025; Lin et al. 2025). Recently, Reinforcement Fine-Tuning (RFT) has been proposed to enhance GUI agents (Lu et al. 2025; Luo et al. 2025), and achieves considerable improvements compared to previous approaches.
Despite the significant momentum of GUI agent techniques, their practical application is hindered by several limitations in both training and inference stages:
  • P1: A Dilemma for Reasoning Designs: An elaborate reasoning process not only degrades grounding accuracy but also increases inference latency, while a “No Thinking” approach exhibits low accuracy for predicting non-grounding action types (Shen et al. 2025).
  • P2: Ineffective Reward: Agents often get stuck on complex interfaces and receive no effective learning signal (i.e., sparse reward). Besides, simple binary feedback (correct/incorrect), a design used by many existing methods (Lu et al. 2025; Luo et al. 2025) may fail to endow agents with the ability to perform precise localization.
@@
  • To overcome P1, UI-AGILE applies a “Simple Thinking” strategy that employs reasoning through thoughts with appropriate lengths. “Simple Thinking” effectively balances the improvement of both the core grounding task and the prediction of non-grounding action types.
  • To tackle P2, UI-AGILE harnesses a design of continuous grounding reward for the RFT stage to incentivize more precise localization to the target’s center. Furthermore, UI-AGILE employs cropping-based resampling to dynamically adjust the difficulty of training samples to avoid ineffective training with zero reward.
@@
Extensive experiments validate the effectiveness of our methods. Trained on about only 9k samples for just 2 epochs, UI-AGILE shows superior performance, while also showcasing strong general agent capabilities. Furthermore, our inference method can act as a plug-and-play enhancement for a wide range of existing agents, improving the accuracy of some existing open-source models.
2 Related Work
Reinforcement Learning (RL) for Large Models
Recently, reinforcement learning (RL) for training large models has gained significant momentum. The focus has shifted from traditional policy optimization algorithms like PPO (Schulman et al. 2017), towards alignment-centric methods such as DPO (Rafailov et al. 2023), and more recent rule-based algorithms like GRPO (Shao et al. 2024). These algorithms have achieved remarkable success in enhancing the reasoning capabilities of large models on complex tasks, with models like OpenAI O1 (Jaech et al. 2024) and DeepSeek-R1 (DeepSeek-AI et al. 2025). The efficacy of these approaches promotes their rapid extension into the multimodal domain (Chen et al. 2025; Chen, Luo, and Li 2025; Liu et al. 2025b; Peng et al. 2025; Meng et al. 2025).
GUI Agents
The field of GUI agents has evolved rapidly (Wang et al. 2024; Li and Huang 2025). Following early works like CogAgent (Hong et al. 2024) and SeeClick (Cheng et al. 2024), most studies rely on SFT to operate directly on visual inputs. Show-UI (Lin et al. 2025) innovates on visual processing efficiency. OS-Atlas (Wu et al. 2025), UGround (Gou et al. 2025) and Aria-UI (Yang et al. 2024) propose novel, large-scale pipelines to collect and synthesize millions of GUI agent trajectories, significantly improving model generalization. Aguvis (Xu et al. 2024) introduced a two-stage training process that explicitly uses VLM-generated Chain-of-Thought (CoT) data to teach planning and reasoning. JEDI (Xie et al. 2025) constructs a refusal part by mismatching existing instructions with unrelated screenshots. Standing out in complexity and scale, UI-TARS (Qin et al. 2025) utilizes the largest dataset and the most intricate training pipeline, which involves SFT and DPO on human annotated CoT data to improve performance. This data-intensive scaling has motivated a shift towards more effective RL paradigms, first explored by UI-R1 (Lu et al. 2025) and GUI-R1 (Luo et al. 2025). InfiGUI-R1 (Liu et al. 2025a) employs Spatial Reasoning Distillation and RL to enhance planning and error recovery capabilities. GUI-G1 (Zhou et al. 2025) leverages Hit-based reward and IoU-based reward for improving GUI agents.
@@ 3 Our Framework UI-AGILE
Fig. 1 provides an overview of UI-AGILE, which aims at improving both training and inference stages of GUI agents.
At training time, the model generates multiple responses from an image and instruction, which are evaluated by reward functions, including “Simple Thinking” reward for efficient reasoning (Sec. 3.1) and continuous grounding reward (Sec. 3.2) for precise localization. If a training sample proves too difficult (i.e., receive a grounding reward of zero for all generated responses), the image will be cropped to simplify the task, and the model resamples new responses from this modified input.
At inference time, UI-AGILE applies decomposed grounding with selection (Sec. 3.4) to enhance grounding on the high-resolution displays common in modern applications, making GUI agents more practical for real-world use.
@@
Lu et al. (2025) posit that excessive reasoning is not essential for GUI grounding and can even be detrimental. However, the complete role of GUI agents extends beyond mere grounding, and deciding next action (e.g., click or type) inherently demands a foundational level of reasoning.
To reconcile the dilemma of whether to apply reasoning for enhancing action-type prediction or discard excessive reasoning to avoid ineffective grounding, we propose a “Simple Thinking” strategy. It encourages thoughts with an appropriate length, operationalized through a specialized reward function. It also reduces training and inference costs, a practical consideration for real-world deployment.
The reward $R_{think}$ for the thought process is defined as:
@@
$R_{\text{length}}(L)=\begin{cases}1.0&\text{if }l_{\text{ideal\_start}}<L\leq l_{\text{ideal\_end}}\\ \\ \frac{1}{2}\left(1-\cos\left(\pi\frac{L-l_{\text{min}}}{l_{\text{ideal\_start}}-l_{\text{min}}}\right)\right)&\text{if }l_{\text{min}}<L\leq l_{\text{ideal\_start}}\\ \\ \frac{1}{2}\left(1+\cos\left(\pi\frac{L-l_{\text{ideal\_end}}}{l_{\text{max}}-l_{\text{ideal\_end}}}\right)\right)&\text{if }l_{\text{ideal\_end}}<L<l_{\text{max}}\\ \\ 0&\text{otherwise}\end{cases}$ (2)
@@
“Simple Thinking” defines an ideal range where the reward is maximized, encouraging thoughts that are neither too brief (“under-thinking”) nor too verbose (“over-thinking”). Outside this ideal range, it uses the cosine function for smooth degradation down to a reward of zero at the absolute bounds. This smooth, non-linear penalty provides a more stable learning signal for RL than a hard cliff. Furthermore, the additional bonus for syntactically complete thoughts discourages incomplete reasoning, thereby ensuring better training stability.
@@ 3.2 Continuous Grounding Reward for Precise Localization (P2)
Unlike general visual grounding tasks which compute the IoU between the predicted bounding box and the ground truth box, GUI agents usually predict a click point and receive a simple binary reward in rule-based RFT (Lu et al. 2025; Luo et al. 2025): a reward of 1 for a correct prediction (e.g., inside the ground truth box) and 0 otherwise.
However, this binary reward is insufficient for high-precision control, as it equally rewards clicks on an element’s edge and at its center. This non-discriminatory feedback misguides the model to learn an element’s boundaries rather than its semantic core.
To resolve this issue, we introduce a continuous grounding reward. Instead of a binary outcome, the continuous reward is calculated as a function of the distance from the predicted point to the center of the ground-truth bounding box:
$R(x,y)=\begin{cases}1+\exp(-4\cdot d_{\text{norm}}^{2})&\text{if }(x,y)\in\text{BBox}\\ 0&\text{otherwise}\end{cases}$ (3)
@@
We employ the Chebyshev distance instead of the Euclidean distance because the reward contours generated by the Chebyshev distance are squares, which geometrically align with the rectangular shape of GUI bounding boxes. This alignment provides a more logical and consistent reward landscape, strongly incentivizing the agent to minimize its deviation along both axes to achieve a high reward.
@@ 3.3 Cropping-Based Resampling for Sparse Reward Mitigation (P2)
During GRPO training, GUI agents often face the sparse reward challenge, particularly on complex tasks. When the model fails to place its prediction within the correct bounding box for a given screenshot over all generations, it receives reward of 0. The lack of positive signals can lead to training stagnation, and difficult samples cannot contribute to model improvement.
Inspired by curriculum learning (Bengio et al. 2009), which presents training examples in an easy-to-hard progression, we propose Cropping-based Resampling. This method acts as a dynamic difficulty adjustment mechanism: if a sample yields zero reward over all generations, we deem it too difficult and reduce its complexity by cropping the screenshot. The resulting smaller crop is guaranteed to fully contain the ground-truth bounding box.
A naive implementation is to center the ground-truth bounding box (bbox) in the new cropping, but it may fail to endow the model with the capability to perform robust localization: the model would learn a trivial shortcut, such as developing a bias for predicting the image center. We opt to employ a scanning approach as illustrated in Alg. 1, ensuring that the cropped image fully contains the ground-truth bounding box. It firstly determines the size of cropping based on a predefined ratio (lines 1-2). Then, the horizontal stride $step_{x}$ is set to the difference between the cropping width and the bounding box width, while the vertical stride $step_{y}$ is set to the difference between their respective heights (lines 3-5). After that, it iterates through all possible cropping windows from left-to-right and top-to-bottom across the original screenshot with the horizontal stride and the vertical stride (lines 7-18). The first window that fully contains the ground-truth bbox is selected as the new, resampled input (lines 12-16). Fig. 2 illustrates how our scanning approach identifies valid cropping windows that fully contain the ground-truth bounding box.
@@ 3.4 Decomposed Grounding with Selection for Visual Noise Reduction (P3)
Modern electronic devices feature high-resolution displays (e.g., 3840x2160), which, when converted into tokens for a VLM, can result in an overwhelmingly long sequence (e.g., over 10,000 tokens). We hypothesize that a significant portion of these tokens represent irrelevant background information, i.e., the visual noise issue P1 mentioned in Sec. 1.
To validate this hypothesis, we conduct a preliminary experiment on ScreenSpot-Pro (Li et al. 2025). We apply the cropping method in Sec. 3.3, but for the purpose of creating a controlled test environment. For each original screenshot, we crop it to 1024x1024, ensuring the ground-truth bounding box is contained within the frame. On this new dataset, the grounding accuracy of UGround-V1-7B (Gou et al. 2025) shows a significant improvement from 31.6 to 56.0, verifying our hypothesis.
@@
[…]
@@
It is straightforward to consider applying the cropping-based method to alleviate the visual noise. However, during inference, the location of the ground-truth bounding box is unknown, making such oracle-based cropping impossible. To overcome this challenge, we propose a multi-stage decomposed grounding with selection method as shown in the right part of Fig. 1. It reduces visual noise by cropping an image into several sub-images and predicts the coordinate individually, while trying to maintain full ground-truth bounding box area. The detailed process is as follows:
@@
    Decomposition: The input screenshot is first divided into several overlapping sub-images, breaking down the high-resolution screen into smaller, more manageable regions.
@@
    Candidate Generation: The GUI agent performs grounding independently on each sub-image and predicts a coordinate for the sub-image. Then, we use the predicted coordinates of sub-images belonging to a screenshot as the candidate coordinates for the screenshot.
@@
Analysis of Inference Cost.
We analyze the inference latency of using decomposed grounding with selection by breaking it down into three primary stages: prefilling, decoding and selection stage.
Counter-intuitively, our approach can theoretically accelerate the computationally-heavy prefilling stage. Recall that the self-attention mechanism has a quadratic time complexity $\mathcal{O}(n^{2})$ concerning the input sequence length $n$ (Vaswani et al. 2017). By splitting a single large image with $n$ tokens into 4 sub-images (each with roughly n/4 tokens), the total computational cost for attention scales proportionally to $4\times(\frac{n}{4})^{2}=\frac{n^{2}}{4}$, suggesting a theoretical speedup. Such cost reduction far outweighs the slightly increased overhead of repeatedly processing the text prompt for each sub-image.
While the decoding process runs for each sub-image, the small number of output tokens (compared to image tokens) per run ensures that the cumulative decoding cost will not increase much.
@@
Overall, the cost of applying decomposed grounding with selection is low. In addition to the above analysis, we provide results on actual running time in Sec. 4.5. Crucially, we believe this overhead could be eliminated or even reversed with future optimizations in inference engines tailored for this “many small requests” workload.

### v4

@@ Abstract
The emergence of Multimodal Large Language Models has driven significant advances in Graphical User Interface (GUI) agent capabilities. Nevertheless, we recognize that existing training and inference techniques for GUI agents still suffer from a dilemma for reasoning designs, ineffective reward, and visual noise. To address these issues, we introduce UI-AGILE for enhancing GUI agents at both training and inference. For training, we propose a suite of improvements to the Reinforcement Fine-Tuning (RFT) process: 1) a “Simple Thinking” strategy to balance planning with speed and grounding accuracy, 2) a continuous reward function to incentivize high-precision grounding, and 3) a cropping-based resampling strategy to mitigate the sparse reward problem and improve learning on complex tasks. For inference, we present decomposed grounding with selection to improve grounding accuracy on high-resolution displays by breaking the image into smaller, manageable parts. Experiments show that UI-AGILE exhibits better grounding performance than state-of-the-art methods on benchmarks ScreenSpot-Pro and ScreenSpot-v2, while showing strong agent capabilities. For instance, using both our training and inference enhancements brings 23% grounding accuracy improvement over the best baseline on ScreenSpot-Pro. The implementation of UI-AGILE is provided at https://github.com/KDEGroup/UI-AGILE.
Key Words: GUI Agent, GUI Grounding, MLLM, RL
@@ 1 Introduction
Driven by the growing capabilities of Multimodal Large Language Models, Graphical User Interface (GUI) agents, which execute tasks by understanding screenshots and user instructions, are advancing rapidly [38].
Prior methods for GUI agents mostly rely on Supervised Fine-Tuning (SFT), requiring a large amount of human-annotated or synthesized data for teaching agents how to plan its actions and grounding [7, 33, 14]. Recently, Reinforcement Fine-Tuning (RFT) has been proposed to enhance GUI agents [17, 18], and achieves considerable improvements compared to previous approaches.
Despite the significant momentum of GUI agent techniques, we recognize that their practical application is hindered by several limitations in both training and inference stages (see Sec. 2 for details):
  • P1: A Dilemma for Reasoning Designs: An elaborate reasoning process not only degrades grounding accuracy but also significantly increases training time and inference latency, while a “No Thinking” approach exhibits low accuracy for predicting non-grounding actions [27].
  • P2: Ineffective Reward: GUI agents often get stuck on complex interfaces and receive no effective learning signal (i.e., sparse reward). Besides, simple binary feedback (correct/incorrect), a design used by many existing methods [17, 18] may fail to endow agents with the ability to perform precise localization.
@@
  • To overcome P1, UI-AGILE applies a “Simple Thinking” strategy that employs reasoning with appropriate lengths when trained on GUI grounding data, while not restricting the reasoning length when trained on data for other actions. “Simple Thinking” effectively reduces training and inference cost while also balancing both the core grounding task and the prediction of non-grounding actions.
  • To tackle P2, UI-AGILE employs cropping-based resampling to dynamically adjust the difficulty of training samples to avoid ineffective training with zero reward. Furthermore, UI-AGILE harnesses a design of continuous grounding reward for the RFT stage to incentivize more precise localization to the target’s center.
@@
Extensive experiments validate the effectiveness of our methods. Trained on about only 9k samples for just 2 epochs, based on Qwen2.5-VL [1], UI-AGILE shows superior grounding performance, while also showcasing strong general agent capabilities. Furthermore, our inference method can act as a plug-and-play enhancement for a wide range of existing agents, improving the accuracy of some open-source models.
2 Preliminary Analysis
To substantiate the problems (P1-P3) outlined in Sec. 1, we conduct a series of pilot studies with the same settings as our experiments in Sec. 4.
2.1 Dilemma of Reasoning (P1)
We investigate the trade-off between reasoning complexity, grounding accuracy, and agent training time. A preliminary analysis, detailed further in our ablation study (Fig. 3 in Sec. 4.4), reveals a clear dilemma:
  • “No Thinking” models (which directly output actions) achieve slightly better grounding accuracy but suffer from significant performance degradation on general agent abilities.
  • “Normal Thinking” models (using elaborate reasoning) achieve better agent abilities but at the cost of degraded grounding accuracy and significantly increased training time (requiring 2x more time than “No Thinking” approach).
The observation confirms that both of the two strategies adopted by existing works have their limitations, motivating us to adopt “Simple Thinking” strategy to balance grounding, planning, and efficiency.
2.2 Ineffective Reward (P2)
To quantify the ineffective reward problem, we analyze the normal training process of the GUI agent with RFT. We track the frequency of ineffective steps, where all responses generated for a training sample receive zero reward, providing no learning signal.
During the first training epoch, only 61.8% of training steps were successful on the first attempt. This implies that, without intervention, 19.1% to 38.2% training samples (here we have 2 training samples per step) will provide no learning signal. The observation highlights the critical need for a mechanism to mitigate sparse rewards, motivating us to design the cropping-based resampling strategy.
2.3 Visual Noise (P3)
Modern electronic devices feature high-resolution displays (e.g., 3840x2160), which, when converted into tokens for a VLM, can result in an overwhelmingly long sequence (e.g., over 10,000 tokens). We hypothesize that a significant portion of these tokens represent irrelevant background information, i.e., visual noise.
To validate this hypothesis, we conduct a preliminary experiment on ScreenSpot-Pro [12]. We apply our cropping method that will be introduced in Sec. 3.3, but for the purpose of creating a controlled test environment. For each original screenshot, we crop it to 1024x1024, ensuring the ground-truth bounding box is contained within the frame. On this new dataset, the grounding accuracy of UGround-V1-7B [7] shows a significant improvement from 31.6 to 56.0, verifying our hypothesis.
@@ 3 Our Framework UI-AGILE
In this section, we will illustrate UI-AGILE. Fig. 1 provides an overview of UI-AGILE, which aims at improving both training and inference stages of GUI agents.
During training, it generates multiple responses from an image and instruction, which are evaluated by reward functions, including “Simple Thinking” reward for efficient reasoning (Sec. 3.1) and continuous grounding reward (Sec. 3.2) for precise localization. If a training sample proves too difficult (i.e., receive a grounding reward of zero for all generated responses), the image will be cropped to simplify the task, and the model resamples new responses from this modified input (Sec. 3.3).
During inference, it applies decomposed grounding with selection (Sec. 3.4) to enhance grounding on the high-resolution displays common in modern applications, making GUI agents more practical for real-world use.
@@
Lu et al. [17] posit that excessive reasoning is not essential for GUI grounding and can even be detrimental. In addition, excessive reasoning can significantly increase training time and inference time. However, the complete role of GUI agents extends beyond mere grounding, and deciding next action (e.g., click or type) inherently demands a foundational level of reasoning.
To reconcile the dilemma of whether to apply reasoning or discard excessive reasoning, we propose “Simple Thinking”. It encourages thoughts with an appropriate length when trained with GUI grounding data, operationalized through a specialized reward function. When trained on data for other actions, it does not restrict thought length.
The reward $R_{think}$ is defined as:
@@
$R_{\text{length}}(L)$ $=\begin{cases}1.0&\text{if }l_{\text{ideal\_start}}<L\leq l_{\text{ideal\_end}}\\ F1&\text{if }l_{\text{min}}<L\leq l_{\text{ideal\_start}}\\ F2&\text{if }l_{\text{ideal\_end}}<L<l_{\text{max}}\\ 0&\text{otherwise}\end{cases}$ (2)
$F1$ $=\frac{1}{2}\left(1-\cos\left(\pi\frac{L-l_{\text{min}}}{l_{\text{ideal\_start}}-l_{\text{min}}}\right)\right)$ (3)
$F2$ $=\frac{1}{2}\left(1+\cos\left(\pi\frac{L-l_{\text{ideal\_end}}}{l_{\text{max}}-l_{\text{ideal\_end}}}\right)\right)$ (4)
@@
This function defines an ideal range where the reward is maximized, encouraging thoughts that are neither too brief (“under-thinking”) nor too verbose (“over-thinking”). Outside this ideal range, it uses the cosine function for smooth degradation down to a reward of zero at the absolute bounds. This smooth, non-linear penalty, inspired by the Hann window [21], a concept fundamental to signal processing, provides a more stable learning signal for RL than a hard cliff. Furthermore, the additional bonus for syntactically complete thoughts discourages incomplete reasoning, thereby ensuring better training stability.
@@ 3.2 Continuous Grounding Reward for Precise Localization (P2)
Unlike general visual grounding tasks which compute the IoU between the predicted bounding box and the ground truth box, GUI agents usually predict a click point and receive a simple binary reward in rule-based RFT [17, 18]: a reward of 1 for a correct prediction (e.g., inside the ground truth box) and 0 otherwise. However, this binary reward is insufficient for high-precision control, as it equally rewards clicks on an element’s edge and at its center. This non-discriminatory feedback misguides the model to learn an element’s boundaries rather than its semantic core.
To resolve this issue, we introduce a continuous grounding reward. It is calculated as a function of the distance from the predicted point to the center of the ground-truth bounding box:
$R(x,y)=\begin{cases}1+\exp(-C\cdot d_{\text{norm}}^{2})&\text{if }(x,y)\in\text{BBox}\\ 0&\text{otherwise}\end{cases}$ (5)
@@
We employ the Chebyshev distance instead of the Euclidean distance because the reward contours generated by the Chebyshev distance are squares, which geometrically align with the rectangular shape of GUI bounding boxes. The exponential term $\exp(-C\cdot d_{\text{norm}}^{2})$ provides a Gaussian-like reward landscape [21] with a peak at the center. This creates a strong, differentiable gradient signal that guides the agent precisely to the element’s semantic core rather than its boundaries.
@@ 3.3 Cropping-Based Resampling for Sparse Reward Mitigation (P2)
During GRPO training, GUI agents often face the sparse reward challenge, particularly on complex tasks. When the model fails to place its prediction within the correct bounding box for a given screenshot over all generations, it receives no positive signals, leading to training stagnation, thus difficult samples cannot contribute to model improvement.
Inspired by curriculum learning [2], which presents training examples in an easy-to-hard progression, we propose Cropping-based Resampling. This method acts as a dynamic difficulty adjustment mechanism: if a sample yields zero reward over all generations, we deem it too difficult and reduce its complexity by cropping the screenshot. The resulting smaller crop is guaranteed to fully contain the ground-truth bounding box.
A naive implementation is to center the ground-truth bounding box (bbox) in the new cropping, but the model would learn a trivial shortcut, such as developing a bias for predicting the image center. We opt to employ a scanning approach as illustrated in Alg. 1, ensuring that the cropped image fully contains the ground-truth bounding box. It firstly determines the size of cropping based on a predefined ratio (lines 1-2). Then, the horizontal stride $step_{x}$ is set to the difference between the cropping width and the bounding box width, while the vertical stride $step_{y}$ is set to the difference between their respective heights (lines 3-5). After that, it iterates through all possible cropping windows from left-to-right and top-to-bottom across the original screenshot with the horizontal stride and the vertical stride (lines 7-18). A random one of windows that fully contains the ground-truth bbox is selected as the new, resampled input (lines 12-16). Fig. 2 illustrates how our scanning approach identifies valid cropping windows that fully contain the ground-truth bounding box.
@@ 3.4 Decomposed Grounding with Selection for Visual Noise Reduction (P3)
1:  $(w_{o},h_{o})\leftarrow GetWidthAndHeight(Image)$
2:  $(w_{crop},h_{crop})\leftarrow(w_{o}\times f,h_{o}\times f)$
3:  $(w_{b},h_{b})\leftarrow(Bbox[2]-Bbox[0],Bbox[3]-Bbox[1])$
4:  $step_{x}\leftarrow w_{crop}-w_{b}$
5:  $step_{y}\leftarrow h_{crop}-h_{b}$
6:  $T\leftarrow\emptyset$
@@
It is straightforward to consider applying the cropping-based method to alleviate the visual noise. However, during inference, the location of the ground-truth bounding box is unknown, making such oracle-based cropping impossible. To solve this issue, we propose a multi-stage decomposed grounding with selection method as shown in the right part of Fig. 1. It reduces visual noise by cropping an image into several sub-images and predicts the coordinate individually, while trying to maintain full ground-truth bounding box area. The detailed process is as follows:
@@
    Decomposition: The input screenshot is divided into several overlapping sub-images, breaking down the high-resolution screen into smaller, manageable regions.
@@
    Candidate Generation: The model performs grounding independently on each sub-image to get coordinates, which serve as candidate points.
@@
Analysis of Inference Cost. We analyze the inference latency of decomposed grounding with selection by breaking it down into prefilling, decoding and selection stages.
Counter-intuitively, our approach can theoretically accelerate the computationally-heavy prefilling stage. Recall that the self-attention mechanism has a quadratic time complexity $\mathcal{O}(n^{2})$ concerning the input sequence length $n$ [28]. By splitting a single large image with $n$ tokens into 4 sub-images (each with roughly n/4 tokens), the total computational cost for attention scales proportionally to $4\times(\frac{n}{4})^{2}=\frac{n^{2}}{4}$, suggesting a theoretical speedup. Such cost reduction far outweighs the slightly increased overhead of repeatedly processing the text prompt for each sub-image.
While the decoding process runs for each sub-image, the small number of output tokens (applying “Simple Thinking”) per run ensures that the cumulative decoding cost will not increase much.
@@
Overall, the cost of applying decomposed grounding with selection is low. In addition to the above analysis, we provide results on actual running time in Appendix D. Crucially, we believe this overhead could be eliminated or even reversed with future optimizations in inference engines tailored for this “many small requests” workload.